9.2.Two point charges in air at a distance of 20 cm from each other


interact with a certain force. At what distance from each other should


these charges be placed in oil to'obtain the same force of interaction?

1
Expert's answer
2022-02-01T11:39:13-0500

"F_{air}= \frac{1}{4\piε_0} \frac{q_1q_2}{r_1}.... ...(i)"

"F_{medium}= \frac{1}{4\piε_0k} \frac{q_1q_2}{d^2}.... ...(ii)"

Substituting equation (i) by (ii) we get,

"\frac{1}{r^2}=\frac{1}{k}×\frac{1}{d^2}"

"d= \\sqrt{\\frac{r^2}{k}}\n \n\n \n\u200b"

r=20cm

k=5

"d= \sqrt{\frac{20^2}{5}} =\sqrt{80} = 8.94cm"

"d=8.94\u00d710^{-2}M"